OpenStudy (anonymous):

(-1, 3) and perpendicular to y = -one thirdx + 7

jigglypuff314 (jigglypuff314):

a perpendicular line has a flipped opposite slope. slope of y = -1/3 x + 7 is -1/3 the flipped of it would be -3 then opposite would be 3 so the line perpendicular to it would have a slope of 3 so given m=3 and a point of (-1, 3) now do you know how to do it? :)

OpenStudy (anonymous):

im doing this test and this test I have no clue about and I try to learn it but I never get it I suck at math :(

jigglypuff314 (jigglypuff314):

no one sucks at math, you just need more practice and proper teaching :) I'll do my best to help in the ways that I can! :D

OpenStudy (anonymous):

y = 3x - 12 y = 3x + 6 y = 3x - 6 y = 3x

OpenStudy (anonymous):

your saying the slope is 3 so it would be d? bc the 1/3 cancel out the 1 out to go to 3

jigglypuff314 (jigglypuff314):

not quite :) slope = 3 means that in y = mx + b would become y = 3x + b then to find b, you can temporarily plug in the given point (-1, 3) as (x, y) so when you plug in -1 for x and 3 for y you get 3 = 3(-1) + b solve for b and tell me what you get :)

OpenStudy (anonymous):

b = -6

jigglypuff314 (jigglypuff314):

close! :) 3 = 3(-1) + b 3 = -3 + b then add 3 to both sides +3 +3 6 = b so b = 6

jigglypuff314 (jigglypuff314):

then you can plug b into y = 3x + b and get... y = 3x + 6 tah-dah! :)

OpenStudy (anonymous):

can you help me with more ?

jigglypuff314 (jigglypuff314):

sure :)

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of one half

OpenStudy (anonymous):

there a lot more :(

OpenStudy (anonymous):

it the same but with the one half

jigglypuff314 (jigglypuff314):

slope = m = 1/2 so for y = mx + b turns into y = 1/2 x + b then you can temporarily plug in the point (1, -4) which mean plug in 1 for x and -4 for y so solve for b when given (-4) = 1/2 (1) + b

OpenStudy (anonymous):

1

OpenStudy (anonymous):

no that wrong

jigglypuff314 (jigglypuff314):

try again then :)

jigglypuff314 (jigglypuff314):

(hint: add 1/2 to both sides)

OpenStudy (anonymous):

so 1/2

OpenStudy (anonymous):

bc it cancel it out right

jigglypuff314 (jigglypuff314):

(-4) = 1/2 (1) + b -4 = 1/2 + b then subtract 1/2 from both sides -1/2 -1/2 -4 1/2 = b so b = - 4 1/2 or -4.5 or -9/2

jigglypuff314 (jigglypuff314):

which you would then plug into y = 1/2 x + b so y = 1/2 x - 4 1/2

OpenStudy (anonymous):

y - 4 = one half(x - 1) y + 4 = 2(x - 1) y + 4 = one half(x - 1) y - 4 = 2(x + 1) but there 4 with all the answers

jigglypuff314 (jigglypuff314):

ooooohhhhhhhhh point slope form: y - y1 = m(x - x1) when m = slope and given the point (x1, y1)

jigglypuff314 (jigglypuff314):

so plug in m = 1/2 x1 = 1 and y1 = -4 into y - y1 = m(x - x1)

jigglypuff314 (jigglypuff314):

no canceling out involved, but plug in :)

OpenStudy (anonymous):

what

OpenStudy (anonymous):

im an little lost

OpenStudy (anonymous):

1,1 1,4 ?

jigglypuff314 (jigglypuff314):

y - y1 = m (x - x1) is the point-slope form for a line, plug in what they gave you (-4) (1/2) (1)

OpenStudy (anonymous):

y - 4 = one half(x - 1)

jigglypuff314 (jigglypuff314):

close! remember that -(-) = + so -(-4) = +4

jigglypuff314 (jigglypuff314):

so y + 4 = one half(x - 1)

OpenStudy (anonymous):

so would the y - be an +

jigglypuff314 (jigglypuff314):

yep! :)

OpenStudy (anonymous):

so c y + 4 = one half(x - 1)

jigglypuff314 (jigglypuff314):

correct! :)

OpenStudy (anonymous):

:) I have 7 left I been doing some while waiting

OpenStudy (anonymous):

would you care to help me again

jigglypuff314 (jigglypuff314):

sure :) although I have about 20 minutes until my bed time :P

OpenStudy (anonymous):

:/ alright I will look for the hardest one

OpenStudy (anonymous):

alright all where hard so I just went to the next one /Choose the equation of the line passing through the point (-2, 6) and parallel to y = one halfx - 8.

jigglypuff314 (jigglypuff314):

parallel means that they have equal slope y = one halfx - 8 ^ 1/2 is the slope so the new line has a slope of 1/2 so so far we have y = 1/2 x + b

OpenStudy (anonymous):

y = 1/2 x + b alright

jigglypuff314 (jigglypuff314):

then temporarily plug in point (-2, 6) so given 6 = (-2)(1/2) + b solve for b

OpenStudy (anonymous):

y = one halfx + 5

jigglypuff314 (jigglypuff314):

close :) 6 = (-2)(1/2) + b 6 = -1 + b so add 1 to both sides so b = 7

OpenStudy (anonymous):

take off the 5 put on 7

jigglypuff314 (jigglypuff314):

yep :)

jigglypuff314 (jigglypuff314):

see, you're starting to get it now! :D

OpenStudy (anonymous):

I guess it time of bed right and yea but last year I fail badly I had an 7% but I try and try and I just really bad am in geometry now and have an b+ but I have to remake this up so I want to learn right not just be stupid and give up in high school

jigglypuff314 (jigglypuff314):

it's alright :) I'm failing math too :P if you ever need help when I'm not around, you can tag people by putting a "@" in front of their username some people that can help are ehuman inkyvoyd esshotwired nikato RadEn ranga shamil98 SolomonZelman

OpenStudy (anonymous):

:) thank you well I hope you have an amazing night while I get more help :P

jigglypuff314 (jigglypuff314):

glad I could help :) good night

OpenStudy (anonymous):

good night